Yellow:

Applications/Percents: Evan spent 20 hours doing homework last week. This week he spent 25 hours doing homework. He says that he spent 125% more time doing homework this week. Is he correct? Show your work to justify your decision.

sak123:

screenshot pls

Yellow:

1 attachment
ZombieLover:

25/20 =1.25 -1 =.25 x 100 =25% more time spent on his HW.

sak123:

he is right, but you will have to try to figure it out

Yellow:

dude i have no idea how

sak123:

125% of 20 =25

Yellow:

Oh

Yellow:

Is that all that needs to be typed?

sak123:

yes ig

Yellow:

Oh ok thenn, thank you loll

sak123:

Yw